Circular No. 7782

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001is IN NGC 1961 Y. L. Qiu and J. Y. Hu, Beijing Astronomical Observatory (BAO), re-

port the discovery by the BAO Supernova Survey of an apparent supernova (mag ∼ 17.6) in unfiltered images taken with the BAO 0.6-m telescope on Dec. 22.56 and 23.73 UT; the new object appeared at mag 17.9 on Dec. 26.53. SN 2001is is located at α = 5^{h}42^{m}09^{s}.07, δ = +69^{o}21′54′′.8 (equinox 2000.0), which is 23′′.2 east and 48′′.2 south of the nucleus of NGC 1961. An image taken with the same telescope on Dec. 14 showed no star at the position of the new object (limiting mag ∼ 18.5). SN 2001is was apparently detected independently by LOTOSS (cf. IAUC 7514), as reported by M. Papenkova (University of California, Berkeley) and M. Schwartz (Cottage Grove, OR), at the following unfiltered CCD magnitudes: 1998 Sept. 17.5, [19.0: (KAIT); 2001 Dec. 24.3, ∼ 18.0 (KAIT); 26.4, ∼ 18.0 (Tenagra III 0.5-m telescope). Schwartz provides position end figures 09^{s}.16, 55′′.3, or 25′′.4 east and 49′′.8 south of the nucleus of NGC 1961.

SUPERNOVA 2001it IN MCG +09-25-15 G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles,

Coddenham, England, of an apparent supernova (mag 17.6) on an unfiltered CCD image (limiting mag 18.5) taken on Dec. 22.196 UT. The new object is located at α = 15^{h}06^{m}09^{s}.40, δ = +53^{o}24′41′′.6 (equinox 2000.0; average of two frames from Dec. 22 and 26), which is 7′′ west and 3′′ north of the nucleus of MCG +09-25-15. A follow-up image taken on Dec. 26.130, obtained in good conditions (limiting mag 19.0), clearly shows the object at mag 17.4. A master image taken on Sept. 16 by Boles does not show SN 2001it to a limit of mag ∼ 18.

SUPERNOVA 2000cl IN NGC 3318 R. A. Stathakis, Anglo-Australian Observatory, reports that he and J.

Stevenson obtained a spectrum (range 300–900 nm) of SN 2000cl (IAUC 7432) on Dec. 18 (570 days after discovery) with the Anglo-Australian Telescope. The spectrum clearly that of a type-IIn supernova (rather than the original classification of type II) and resembles that of SN 1988Z at a similar epoch (Turatto et al. 1993, MNRAS 262, 128). The night was cloudy, but the object appeared to be of similar brightness to that at discovery (mag 14.8) via the unfiltered CCD. Further observations are encouraged, especially at other wavelengths.

2001 December 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green